# The POKEMON Speckle Survey of Nearby M Dwarfs. II. Observations of 1125 Targets

### v6

@@ 3.1 Detected Companions
  • In some cases, in particular for the fainter objects, there may be a ${}^{\circ}$ ambiguity in the listed position angles due to the limitations of the reconstructed images described in Section 2.3.
@@ 3.2 Astrometric Precision
To characterize our astrometric precision, we use the fact that DSSI and NESSI observe simultaneously at two wavelengths. This allow us to calculate the residuals in our astrometry by computing the difference between the observed angular separation and position angle in each channel. The angular separation residuals have an average value of 3.1 mas, with a standard deviation of 12 mas. The position angle residuals have an average value of ${}^{\circ}$, with a standard deviation of ${}^{\circ}$. The residuals result from the subtraction of two independent measurements with presumably the same uncertainty, so the subtraction has an uncertainty that is $\sqrt{2}$ larger than the uncertainty of either individual measure. This means that the uncertainty in a single angular separation measure is given by 12 divided by $\sqrt{2}$, or 8.4 mas. The average uncertainty in position angle is then ${}^{\circ}$. These values are larger than those derived in Horch et al. (2017), Colton et al. (2021), or Horch et al. (2021), which also used DSSI and NESSI; this is likely due to the faintless of our targets, and the filterless observations that took place during the 2017 April observing run. In general, filterless observations reduce the precision of our astrometry, but allow us to observe fainter companions.
@@ 3.3 Photometric Precision
[…]
@@ 5 Conclusions and Future Work
This work presents results from the European Space Agency (ESA) space mission. Gaia data are being processed by the Gaia Data Processing and Analysis Consortium (DPAC). Funding for the DPAC is provided by national institutions, in particular the institutions participating in the Gaia MultiLateral Agreement (MLA). The Gaia mission website is https://www.cosmos.esa.int/gaia. The Gaia archive website is https://archives.esac.esa.int/gaia.

### v8

@@ Abstract
journal: AJfacilities: LDT(DSSI), WIYN(NESSI)software: Astropy (Astropy Collaboration et al., 2013), Astroquery (Ginsburg et al., 2019), IPython (Pérez & Granger, 2007), Matplotlib (Hunter, 2007), NumPy (Harris et al., 2020), Pandas (McKinney et al., 2010), SciPy (Virtanen et al., 2020)
@@ 3.1 Detected Companions
  • In some cases, in particular for the fainter objects, there may be a 180 ambiguity in the listed position angles due to the limitations of the reconstructed images described in Section 2.3.
@@ 3.2 Astrometric Precision
To characterize our astrometric precision, we use the fact that DSSI and NESSI observe simultaneously at two wavelengths. This allow us to calculate the residuals in our astrometry by computing the difference between the observed angular separation and position angle in each channel. The angular separation residuals have an average value of 3.1 mas, with a standard deviation of 12 mas. The position angle residuals have an average value of 0.10, with a standard deviation of 1.1. The residuals result from the subtraction of two independent measurements with presumably the same uncertainty, so the subtraction has an uncertainty that is $\sqrt{2}$ larger than the uncertainty of either individual measure. This means that the uncertainty in a single angular separation measure is given by 12 divided by $\sqrt{2}$, or 8.4 mas. The average uncertainty in position angle is then 0.77. These values are larger than those derived in Horch et al. (2017), Colton et al. (2021), or Horch et al. (2021), which also used DSSI and NESSI; this is likely due to the faintless of our targets, and the filterless observations that took place during the 2017 April observing run. In general, filterless observations reduce the precision of our astrometry, but allow us to observe fainter companions.
@@ 3.3 Photometric Precision
We also plot Gaia magnitude difference versus speckle magnitude difference for the 15 observations where seeing times separation is $<0.6$ arcsec2. These observations should be relatively unaffected by non-isoplanicity; the speckle magnitude differences should therefore correlate well with the Gaia magnitude difference values. The correlation is fairly good, although it does appear that there is a systematic offset from an equivalent relation. The residuals from these magnitude differences have an average value of 0.25 mag, with a standard deviation of 0.54 mag. We divide by $\sqrt{2}$, so the average uncertainty in magnitude difference is 0.38.
Figure 6: Left: Difference between speckle magnitude difference and Gaia magnitude difference as function of seeing times separation. This difference between the speckle and Gaia magnitude differences should be $\sim 0$ for small values of seeing times separation, and should increase as seeing times separation increases. This is the trend we see in our data. Right: Gaia magnitude difference versus speckle magnitude difference for the observations where seeing times separation is $<0.6$ arcsec2. The correlation is fairly good, although it does appear that there is a systematic offset from an equivalent relation (dashed orange line).
@@ 5 Conclusions and Future Work
This work presents results from the European Space Agency (ESA) space mission Gaia. Gaia data are being processed by the Gaia Data Processing and Analysis Consortium (DPAC). Funding for the DPAC is provided by national institutions, in particular the institutions participating in the Gaia MultiLateral Agreement (MLA). The Gaia mission website is https://www.cosmos.esa.int/gaia. The Gaia archive website is https://archives.esac.esa.int/gaia.